OpenStudy (anonymous):

Medal for help Parallel lines are cut by a transversal such that the alternate interior angles have measures of 3x + 17 and x + 53 degrees. The value of x is 9 18 35 71

OpenStudy (anonymous):

18,

OpenStudy (anonymous):

\[3x +17 = x +53\]\[3x = x + 36 \]\[2x = 36\]

OpenStudy (anonymous):

x = 18

OpenStudy (anonymous):

thnx:)